Under conditions where ketone bodies are being produced in the liver, how many ATPs can be produced from a molecule of lauric acid if all resulting molecules of acetyl-CoA are converted into β-hydroxybutyrate?

Express your answer using one decimal place.

Net yield of ATP???
Answer #1

Hi! Welcome to HomeworkLib

To convert lauric acid to beta hydroxy butyric acid we have to first do beta hydrolysis of lauric acid to form acetyl CoA and then Acetyl CoA will enter the ketogenetic pathway to form beta hydroxy butyric acid.

So complete beta hydrolysis of lauric acid will produce 6 Acetyl CoA,5NADH2 and 5FADH2.

Now In ketogenesis 2 molecules of Acetyl CoA give one molecule of beta hydroxy butyric acid with use of one molecule of NADH2.

So 6 acetyl CoA will give 3 beta hydroxy butyric acid with use of three NADH2.

Now total production

5NADH2-3NADH2=2NADH2

and 5FADH2.

Calculating ATP equivalents ,1 NADH2 Equals 2.5 ATP and 1 FADH2 equal 1.5 ATP.

TOTAL ATP=(2×2.5)+(1.5×5)=5+7.5=12.5 ATP.